OpenStudy (anonymous):

*Will medal and fan* ---> should this be a or c? If a line has an equation y= (-3/2)x+4, find the slope of a line perpendicular to it. Select one: a. 2/3 b. 3/2 c. -2/3 d. -3/2

OpenStudy (trojanpoem):

dy/dt = -3/2 perpendicular = 2/3

OpenStudy (anonymous):

Thank you! @TrojanPoem

OpenStudy (trojanpoem):

Anytime.

OpenStudy (anonymous):

@ TrojanPoem I NEED HELP ASAP